Paradise Corp. has determined a standard labor cost per unit of $16 (0.50 hour × $32 per hour). Last month, Paradise incurred 958 direct labor hours for which it paid $23,471. The company produced and sold 2,150 units during the month. Calculate the direct labor rate, efficiency, and spending variances. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Round your intermediate calculations to 2 decimal places.)

Direct Labor Rate Variance? Favorable or unfavorable?

Direct Labor Efficiency Variance? Favorable or unfavorable?

Total Direct Labor Spending Variance? Favorable or unfavorable?
Answer #1

Labor rate variance = (standard rate - actual rate)*actual hours

= (32-23471/958)*958 = $7185 favorable

Labor efficiency variance = (standard hours-actual hours)*standard rate

= (2150*0.50-958)*32 = $3744 favorable

Total labor spending variance = 2150*16-23471 = 10929 favorable